OpenStudy (anonymous):

Identify the solution of the compound inequality −3y < 12 and y − 2 ≤ 4.

OpenStudy (anonymous):

okies :D thankss

OpenStudy (anonymous):

no its ok can you help :D?

OpenStudy (anonymous):

bhahahah okies :D

OpenStudy (anonymous):

um. ok. :/

OpenStudy (kittiwitti1):

Hey, I'm back xD I just meant don't wait for me but it seems you didn't want anyone else to help ; ~; I feel so honored cx So yeah... we don't need system of equations, we need substitution after solving for x in either one of the problems. Sorry. D: \[-3y < 12\]Divide both sides by -3. Because the value divided is negative, flip the sign to >.\[Now~we~get~y >-4.\] Then the next one: \[y - 2 \le 4\]\Take off the 2 by adding 2 to both sides.\[y \le 6\]\[So~ -4 < y \le 6.\]

OpenStudy (kittiwitti1):

*correction: we don't use either system of equations or substitution, we just slap the answers into one xD